OpenStudy (anonymous):

You invested $8000 in stocks and bonds, paying 6% and 9% annual interest, respectively. If the total interest earned for the year was $540, how much was invested in stocks and how much in bonds?

OpenStudy (whpalmer4):

okay, we need to write two equations here. one shows how much we've got invested in each class, and the other how much money we made. let S be the amount invested in stocks. let B be the amount invested in bonds. can you write an equation showing the relationship between B, S and 8000?

OpenStudy (anonymous):

actually...yes :O) .06x + .09(8000-x) =540

OpenStudy (anonymous):

now that i look at it, it's EXACTLY the same formula we worked before

OpenStudy (anonymous):

so I came up with 6000 at .06 and 2000 at .09

OpenStudy (whpalmer4):

okay, let's check the result. 6000 * 0.06 is the stock interest = 360 2000 * 0.09 is the bond interest = 180 total interest is 540 = 360+180 looks like a valid solution, good job!

OpenStudy (anonymous):

mind if I hit you with a couple more problems? This class I am taking is required for my degree and unfortunately, is online only. The teacher is on a different campus than the one I attend so I don't get to attend lecture/interact personally with the instructor

OpenStudy (whpalmer4):

if we'd done the two equation route, it might look like S + B = 8000 0.06S + 0.09B = 540 same solution, of course.

OpenStudy (whpalmer4):

sure, fire away

OpenStudy (anonymous):

Two gasoline distributors, A and B, are 228 miles apart on Interstate 20. Distributors A charges $3.02/gal and B charges $2.93/gal. Each charges .2 cents/gal per mile for delivery. Where on Interstate 20 is the cost to the customer the same? Once again, I need help setting it up

OpenStudy (anonymous):

This question is instructor generated, so I am assuming that it isn't possible to charge .2 cents per gal per mile

OpenStudy (whpalmer4):

well, let's imagine it looks like this: A ------------C----------B we can call the distance between A and C (C = customer) x, and then the other chunk is 228-x the customer buys the same amount of gas regardless of who she buys from, right? let's assume that to be true. so the price for customer to buy one gallon from A is distance from A * delivery charge rate for A + cost of one gallon from A price for customer to buy one gallon from B is distance from B * delivery charge rate for B + cost of one gallon from B sound good so far?

OpenStudy (whpalmer4):

if you agree, then just write out those two expressions and set them equal to each other; the price is the same at the spot where the customer wants to be

OpenStudy (anonymous):

*eyes glaze over*

OpenStudy (whpalmer4):

sorry :-) okay, the customer is in the middle, right? let's say it is x miles from A to the customer. from B to the customer it must be 228-x miles, yes?

OpenStudy (anonymous):

it's me being ADHD. I tend to overthink the simple stuff and "get" complex things

OpenStudy (whpalmer4):

okay, well, we can make this simple or complicated, your choice :-)

OpenStudy (anonymous):

right now, i'm trying to take the knowns and plug them into the variables in the correct positions.

OpenStudy (whpalmer4):

so yes, the customer is in the middle, x miles from A and 228-x miles from B. cost of a gallon of gas is the cost for the gas plus the delivery charge. the delivery charge is 0.002 dollars * the distance from the distributor. a gallon bought from A costs $3.02 + x*$0.002 a gallon bought from B costs $2.93 + (228-x)*$0.002

OpenStudy (anonymous):

geez. i'm going to have to start paying you by the hour, hahaha.

OpenStudy (whpalmer4):

we just set 3.02 + 0.002x = 2.93 + 0.002(228-x)

OpenStudy (anonymous):

okay give me a moment to pencil this out

OpenStudy (whpalmer4):

sure!

OpenStudy (whpalmer4):

I'd suggest multiplying both sides by 1000 as a first step to get rid of the pesky decimals

OpenStudy (whpalmer4):

3020 + 2x = 2930 + 2(228-x)

OpenStudy (anonymous):

if i am working this through correctly, i should currently see 90 = 465 - 2x? x = 187.5?

OpenStudy (anonymous):

but if we bumped each side by 1k, do we in the end need to reduce it by 1k?

OpenStudy (whpalmer4):

3020 + 2x = 2930 + 2(228-x) 3020 + 2x = 3386 - 2x 4x = 3386 - 3020 4x = 366 x = 91.5

OpenStudy (whpalmer4):

no, you can add, subtract, multiply, divide, take square roots, logs, you name it, so long as you do the same thing to both sides of the equation, it won't change the solution.

OpenStudy (whpalmer4):

if you have 2 dollar bills and I give you 3 more dollar bills you have 5 dollar bills, right? if we then discover that we didn't have our glasses on and they were actually 10 dollar bills, you still have 5 of them, right?

OpenStudy (anonymous):

lol right

OpenStudy (anonymous):

one thing gets me: why don't we apply the changes to what is enclosed in the ( ) ?

OpenStudy (whpalmer4):

oh, that's easy — we applied it to the outside. a*b*1000 = a*1000 * b, right?

OpenStudy (anonymous):

right

OpenStudy (whpalmer4):

here, we can expand it first, then apply the *1000: 3.02 + 0.002x = 2.93 + 0.002(228-x) 3.02 + 0.002x = 2.93 + 0.002*228 - 0.002*x 3020 + 2x = 2930 + 2*228 - 2*x 3020 + 2x = 3386 - 2x same as we had...

OpenStudy (whpalmer4):

but a bit more work this way

OpenStudy (anonymous):

I'd rather the shorter path. less opportunity to miscalculate.

OpenStudy (whpalmer4):

so let's check our answer. 91.5 miles from A, a gallon of gas costs $3.02 + $0.002*91.5 = $3.203 228-91.5 = 136.5 miles from B, a gallon of gas costs $2.93 + $0.002*136.5 = $3.203

OpenStudy (whpalmer4):

this makes sense — if it costs 2/10 of a cent to carry the gas a mile, then every 10 miles the price changes by 2 cents. we need to have a difference of $3.02-$2.93 = 9 cents, so that's 4.5*10 = 45 miles. if we take 228-45 and then divide the result in half, that's the distance to one of the endpoints.

OpenStudy (anonymous):

This should be simple enough to do in my sleep. Why does this particular setup seem to stonewall me? Where is the difficult part? What am I overlooking?

OpenStudy (whpalmer4):

228-45 = 183, 183/2 = 91.5 :-)

OpenStudy (whpalmer4):

actually, I had to think about it for a minute or two before I wrote down the right thing. don't beat yourself up too much :-)

OpenStudy (anonymous):

where'd you get the 4.5 from the 9 cents?

OpenStudy (anonymous):

other than 4.5 being 9/2

OpenStudy (whpalmer4):

difference of 9 cents per gallon in the underlying price of gas, right? every 10 miles the purchase location changes = 10 * 0.002 cents per mile = 2 cents change in the charge to deliver the gas. if we want a 9 cent change, and we know the distance to change the price by 2 cents, then 9/2 = 4.5 is the number of those distances end to end to change it by 9 cents.

OpenStudy (anonymous):

right on

OpenStudy (whpalmer4):

5 miles changes it by 1 cent, I should have done that to cause less confusion, but I like multiplying by 10 better than I do multiplying by 5, especially with small numbers like this. so 5 miles change on the distance to deliver the gas changes the delivery charge by 1 cent. to overcome the 9 cent per gallon difference in price, we need to move the delivery point by 9 cents/ (1 cent/5 miles) = 45 miles * cents / cents = 45 miles.

OpenStudy (anonymous):

sorry - while working this problem, i was pencilling another. gimme sec to submit answer and see if i got it right

OpenStudy (anonymous):

you do math like i do. you round to the nearest manageable 5 or ten then +-/* as is needed

OpenStudy (whpalmer4):

what I did was 9 cents / (2 cents/10 miles) = 90 miles * cents / 2 cents = 45 miles.

OpenStudy (whpalmer4):

next problem? :-)

OpenStudy (anonymous):

and a tpath to boot.

OpenStudy (anonymous):

jussasec.

OpenStudy (anonymous):

In a new development, 57 one and two bedroom condominiums are sold. Each 1-bedroom condo sold for $130,000 and each 2-bedroom condo sold for $150,000. If sales totaled $8,250,000, how many of each type was sold? we're back with our old friend equation

OpenStudy (whpalmer4):

yep. set it up and knock it down :-)

OpenStudy (anonymous):

130,000x + 150,000(57-x) = 8250,000

OpenStudy (anonymous):

gah i want to knock off a few zeroes.

OpenStudy (whpalmer4):

here's how you could do it in your head... assume all 57 sales are 1 bedroom. 57*130000 = 7410000. that's shy of what we need by 8250000-7410000 = 840000. each 2 bedroom condo instead of a 1 bedroom condo nets us an extra 150000-130000 = 20000. we need 840000/20000=42 condos to be 2 bedroom rather than 1.

OpenStudy (whpalmer4):

if we did this in thousands of dollars instead of dollars, it would be a bit more tractable in the old skull cavity, to be sure

OpenStudy (anonymous):

15 1br/42 2br?

OpenStudy (whpalmer4):

yep. 15+42 = 57, check 15*130 + 42*150 = 1500 + 450 + 6000 + 300 = 7500 + 750 = 8250 tack on 3 zeros to get the original scale back

OpenStudy (anonymous):

You're a lifesaver. The only way I know to repay your kindness is to help others on here.

OpenStudy (whpalmer4):

that would be just fine with me :-)

OpenStudy (anonymous):

hah yeah :) I laid hands on a router years ago and fell head over heels in love

OpenStudy (anonymous):

just need to get used to working with ipv6 and subnetting in v6

OpenStudy (anonymous):

dear God, I've hit the jackpot :)

OpenStudy (anonymous):

now i just need to get more comfortable with VLSM. i can do the addressing while in class, send me home with homework and i falter

OpenStudy (anonymous):

v6 would require a lecture since it's in hex rather than binary

OpenStudy (anonymous):

meaning that i would have to pull out lecture notes to give you the skinny on it

OpenStudy (anonymous):

for now, though, most of the world is still v4, so we're not facing any immediate issues with it, other than cisco just pushed it down into core cisco curriculum last semester

OpenStudy (anonymous):

if you started there in 88, you had to be 20-ish (give or take) so that makes you about my age

OpenStudy (anonymous):

lol i'm 40 going on peter pan

OpenStudy (anonymous):

makes me wonder if you ever met some of the guys from ARPA

OpenStudy (anonymous):

Thankee-sai! I will most definitely look you up again. I really enjoy being shown how to think it through. One last question

OpenStudy (whpalmer4):

sure. make it a doozie :-)

OpenStudy (anonymous):

What is it about this formula I have been using throughout this assignment is so important that my prof is drilling us on it?

OpenStudy (whpalmer4):

oh, it's just the process of recognizing how to write the equation(s) describing the relationships between the various items.

OpenStudy (anonymous):

but they are all ax + b(c-x) = d

OpenStudy (whpalmer4):

and then of course to solve them, but right now it sounds like you're doing it with just one variable, so that's not such a big deal. say I have 85 cents in my pocket, nickels and dimes only, 10 coins. how many of each do I have? the way you're being taught, you would choose either nickels or dimes as x, and the other would be 10-x. then you have 5x + 10(10-x) = 85, 3 nickels and 7 dimes

OpenStudy (anonymous):

holy pellet. you're right. i DID do that one in my head

OpenStudy (whpalmer4):

but that doesn't work so well if you have more than 2 quantities. here's another way you can write it where you can expand to more than 2 quantities without trouble: let n be the number of nickels and d the number of dimes. n + d = 10 5n + 10d = 85 now we have a system of two equations in two unknowns (n and d). we can solve this by substitution or elimination or with matrices. you can see that if we had a problem where I had pennies in the mix as well, it wouldn't be possible to write it with just one variable, right?

OpenStudy (anonymous):

pellet hahahahaha! it changed my expletive into a funny

OpenStudy (whpalmer4):

yeah, keeping us safe from ourselves :-)

OpenStudy (anonymous):

yeeeahhhhh....i'm 40, not 4. I ha.....meh. not worth the diatribe

OpenStudy (whpalmer4):

so I'll solve that system first by substitution: n + d = 10 5n + 10d = 85 well, we pick one of the equations and solve it for one variable. here it's easy: we'll use the first one, solve it for n: n+d = 10 n = 10-d now we substitute that in the other eq, giving us an eq in just one variable, which we solve. then we get the value of the other variable from that and either equation (or most likely, the substitution eq). n = 10-d 5n + 10d = 85 5(10-d) + 10d = 85 50 - 5d + 10d = 85 50 + 5d = 85 5d = 35 d = 7 n = 10-d = 10 -7 = 3 that should look a lot like what you do with one variable, right?

OpenStudy (anonymous):

more or less, yeah. so how'd you get started coming here to tutor?

OpenStudy (whpalmer4):

now elimination: n + d = 10 5n + 10d = 85 here we try to find some combination of multipliers for one or both equations so that we can get the coefficients of one of the variables to be equal but opposite in sign. then we can add the two equations together, giving us an equation in one variable. here I'll multiply the first one by -5: -5n -5d = -50 5n + 10d = 85 -------------- 0n + 5d = 35 5d = 35 d = 7 now plug that back into any of the equations and solve for the other variable n+d = 10 n +7 = 10 n = 3 this approach can be extended to do 3 equations in 3 unknowns, 4 in 4, etc.

OpenStudy (anonymous):

i've got a memory leak somewhere. too much lag in my typing

OpenStudy (whpalmer4):

I don't remember how I stumbled across OpenStudy. it was a few years ago. I'm sort on and off. worked my way up to 99 this summer, then went off on a trip, and after my return didn't spend any time here until just a few weeks ago, I think.

OpenStudy (whpalmer4):

it's fun getting people who think they can't do this stuff to realize that they can, and it's actually fun, not complete tedium. it's even more fun helping people who are really enthusiastic about learning, and I learn new things here just about every day.

OpenStudy (anonymous):

I learned to love math WAAAAAAYYYYYY back when I was in 3rd grade and they taught us "new" math.

OpenStudy (whpalmer4):

unfortunately, there are also a lot of kids who just want answers, don't want to work at all...

OpenStudy (whpalmer4):

do you know the wonderful song by that name by Tom Lehrer?

OpenStudy (anonymous):

pfft it ain't worth it if you don't work for it. don't learn a damn thing

OpenStudy (anonymous):

I DO! i laughed my bum off watching it the first time

OpenStudy (whpalmer4):

my 10 year old son was infected by the Tom Lehrer bug years ago. I keep expected to get a call from the school chiding me about the inappropriate songs he's singing, because he's memorized quite a bit :-)

OpenStudy (anonymous):

fortunately for me, my own two cubs are grown. they can handle their own inappropriate behavior :P

OpenStudy (whpalmer4):

what really piqued my interest was seeing Donald in MathMagic Land http://www.youtube.com/watch?v=AJgkaU08VvY

OpenStudy (whpalmer4):

it was the highlight of each school year :-)